whats the fastest way 2 do division
 

Grade:7

4 Answers

Latika Leekha
8 years ago
Hello student,
If you find the long division method a bit tedious and time consuming you can opt for some other ways like the use of distributive property in order to divide numbers. But, in that case you need to be well versed with the properties as well as have a good speed.
Eg: You can either split the dividend or the divisor to make the problem simpler:
Suppose you have to divide 57 by 3. then instead of carrying out the division directly, you can split teh dividend i.e. 57 into sum of numbers which are both divisible by 3.
i.e. we can write 57= 27+30 and here both 24 and 30 can be easily divided by 3.
so, 57/3 = (27+30)/3
= (27/3) + (30/3)
= 9+10
= 19.
Or you can aslo in some problems split the divisor.. like suppose you need to divide 324/6.
Then you can write 6 = 2*3.
so, 324/6 = 324/(2*3)
Now, 3 and 24 are both divisible by 3 and so we can see 324/3 = 108 and now 108 can be simply divided by 2 to get 54.
Except these, in rpoblems where you are given any three out of dividend, divisor, quotient or remainder, you can use the formula,
dividend = divisor . quotient + remainder.
grenade
8 years ago
u could use the simple algebric identities for the solving of questions
V. Harshit -
8 years ago
u should use  the properties like associative and distributive properties that u have learned .
Kalyani Jayachandran Menon
8 years ago
You need to know the multiplication tables very thoroughly.. With practice, you’ll be able to divide seemingly weird numbers quite easily. For example, you won’t even need to think about problems like 69/23.